OpenStudy (doc.brown):

4/3pi*y^3 = 4/3pi*x^3 + z Simplify to find y

OpenStudy (anonymous):

\[\frac{4\pi y^3}{3}=\frac{4\pi x^3}{3}+z\] What do you think you would do? your variable is already isolated, just some constants with it

OpenStudy (doc.brown):

Is this right?\[\frac43\pi y^3=\frac43\pi x^3 + z\]\[y^3=x^3 + \frac{3z}{4\pi}\]\[y=\sqrt[3]{x^3 + \frac{3z}{4\pi}}\]

OpenStudy (anonymous):

excellent